For any two integers x and y, which of the following suggests that operation of addition is commutative ? 
(A) x + y = y + x (B) x + y > x (C) x – y = y – x (D) x × y = y × x

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Option (A) is right here.
Let us assume a and b are the two integers,
Then, commutative law of addition =
a + b = b + a
Where, a = x
b = y

Therefore, commutative law of addition =
x + y = y + x


Thanks